### v4

# GRAMMAR: Grounded and Modular Methodology for Assessment of Domain-Specific Retrieval-Augmented Language Models

Abstract
Retrieval-augmented Generation (RAG) systems have been actively studied and deployed across various industries to query on domain-specific knowledge base. However, evaluating these systems presents unique challenges due to the scarcity of domain-specific queries and corresponding ground truths, as well as a lack of systematic approaches to diagnosing the cause of failure cases—whether they stem from knowledge deficits or issues related to system robustness. To address these challenges, we introduce GRAMMAR (GRounded And Modular Methodology for Assessment of RAG), an evaluation framework comprising two key elements: 1) a data generation process that leverages relational databases and LLMs to efficiently produce scalable query-answer pairs. This method facilitates the separation of query logic from linguistic variations for enhanced debugging capabilities; and 2) an evaluation framework that differentiates knowledge gaps from robustness and enables the identification of defective modules. Our empirical results underscore the limitations of current reference-free evaluation approaches and the reliability of GRAMMAR to accurately identify model vulnerabilities.
GRAMMAR: Grounded and Modular Methodology for Assessment of Domain-Specific Retrieval-Augmented Language Models
Figure 1: The dual-phase evaluation approach for the GRAMMAR model. The left section outlines the data generation process for creating sets of hypothetically robust ($\mathcal{D}_{\text{robust}}$) and non-robust ($\mathcal{D}_{\text{non-robust}}$) data to test model hypotheses. The right section details the methodology for identifying defective modules. This grouping step accumulates queries with identical semantics and delineates three group types for identifying detective modules. Gap Groups indicate a complete lack of correct responses, reflecting deficiencies in the model’s knowledge or capabilities. Robust Groups show the model’s comprehensive understanding, with correct responses to all queries in the subset. Non-robust Groups highlight the robustness issue, where the model correctly answers some but not all text forms with identical query logic.
1 Introduction
The emergent ability of Large Language Models (LLMs) has advanced a active research and deployment of Retrieval-Augmented Generation (RAG) systems for domain-specific question answering. This trend underscores the critical need for reliable evaluation of these systems, a task that remains intricate and largely unaddressed.
One of the foremost challenges is the scarcity of domain-specific data. Typically, evaluators resort to using data designed for human assessment, as referenced in studies such as Santurkar et al. (2023); Wang et al. (2022); Zhong et al. (2022, 2023); Hendrycks et al. (2021); Choi et al. (2023), which mainly focuses on evaluating commonsense and broad world knowledge. However, the assessment of industrial RAG systems demands knowledge specific to particular domains, such as details about company projects and employees. While user queries can be gathered at the time of system deployment, acquiring accurate ground truths poses a significant challenge. Recent studies have explored the use of reference-free LLM evaluators in the absence of ground truths (Chern et al., 2023; Min et al., 2023; Es et al., 2023). Nonetheless, the reliability of such reference-free evaluation methods remains questionable. Our analysis of two representative methods convincingly demonstrates their false assessment across accurate and inaccurate predictions: only RAGAS demonstrates reliability in identifying correct predictions. In this paper, an automatic data generation methodology is proposed to ensure the accessibility of ground truths (Contribution 1). The methodology leverages relational databases and LLMs to enable the extraction of ground truths via SQL queries.
A second significant challenge relates to the robustness of models (Contribution 2). Specifically, we observe inconsistencies in the models’ responses to queries that, while semantically equivalent, differ in their linguistic presentation, as highlighted in Shen et al. (2023). Such inconsistencies pose a direct challenge to the crucial requirement for consistent performance. Researchers often hypothesize about the factors influencing the robustness or susceptibility of retrieval and language models to certain attributes, questioning what impedes these models from responding accurately to queries. However, the current body of research, including works on adversarial robustness (Alzantot et al., 2018; Li et al., 2020, 2019), lacks the mechanisms to effectively test these hypotheses. To address this gap, we introduce a evaluation workflow, as domonstrated in Figure 1: 1) utilizing the proposed data generation process that creates hypothetically robust data $\mathcal{D}_{\text{robust}}$ and non-robust data $\mathcal{D}_{\text{non-robust}}$ according to a proposed hypothesis, 2) identifying defective modules based on a grouping mechanism, contextual sufficiency analysis and a reliable robustness metric that minimizes knowledge gap interference. Our empirical study highlights the reliability of GRAMMAR to verify the hypothesis and further identify the defective modules.
2 Background
Retrieval-augmented Generation
RAG systems augment LLMs with retrieval for domain-specific use where extensive databases can potentially provide necessary information. A retrieval can be used for retrieving relevant passages or documents. A dense retrieval (Mialon et al., 2023; Lewis et al., 2020; Borgeaud et al., 2022) consists of an embedding model to produce a vector $v_{q}$ for any given query $q$. This vector is then used to identify pertinent text segments normally via dot product calculations with the vectors of text chunks. The most relevant chunks are aggregated to form an additional context $c$, which serves as the factual basis for the response generation, until adding another chunk would surpass the maximum allowable context length. Subsequently, an LLM will process this aggregated context alongside the original query to generate a predicted answer $\hat{a}$ for $q$. In contrast, sparse retrieval computes similarity without parametric embedding models. Keyword matching . For example, sparse retrieval techniques operate without the need for parametric embedding models. For instance, keyword-matching approaches (Manning, 2009) select documents that share the highest number of common words for context retrieval. Additionally, methods like TF-IDF (Chen et al., 2017) factor in the inverse document frequency, emphasizing the significance of less common words which are likely to be more indicative of pertinent content.
Knowledge-intensive Evaluation
Many evaluation protocols of RAG and LLMs focus on open-domain knowledge. Both the queries $Q$ and their ground-truth answers $A$ in the open domains can be easily accessed via online forums, public surveys (Santurkar et al., 2023) and qualification and admission exams that were originally designed for humans Wang et al. (2022); Zhong et al. (2022, 2023); Hendrycks et al. (2021); Choi et al. (2023). Since LLMs are originally trained on open-domain knowledge, these evaluation protocols on RAG focuses more on whether LLMs’ factuality can be improved by retrieval. However, this paper focuses on industrial scenarios where retrieval mostly provides new knowledge to LLMs and ground-truth answers are hardly available. The reference-free evaluation protocol is commonly proposed for this situation, e.g., ARES (Saad-Falcon et al., 2023) and RAGAS (Es et al., 2023). For example, instead of the correctness of answers against the users’ queries, Saad-Falcon et al. (2023) assess answer faithfulness against the context from retrieval modules by profiling LLMs to generate evaluation scores. The results can be skeptical due to the reliability of retrieval and the fairness of LLMs as evaluators. Reference-free evaluation is also utilized to assess LLMs’ open-domain knowledge (Kadavath et al., 2022). SelfCheck (Manakul et al., 2023) relies on the stochastic nature of LLMs for validations, while FactScore (Min et al., 2023) and FacTool (Chern et al., 2023) use Retrieval+LLM as a validation system. FactScore and FactTool are naturally not suitable for evaluating RAG systems that already include retrieval as an inherent component.
3 GRAMMAR: Grounded and Modular Methodology for Assessment of RAG
GRAMMAR is structured around two key components: 1) Grounded Data Generation Method: Grounded Data Generation Method: This approach generates text queries and corresponding answers from an existing knowledge base (relational database) for targeted evaluations. Using SQL as an intermediate representation, it allows for the controlled creation of various text queries with consistent underlying logic. The process and its implementation are detailed in § 3.1 and 3.2, and illustrated in Figure 2. 2) Modular Evaluation Protocol: This protocol assesses robustness by not only identifying proposed vulnerabilities but also pinpointing which component exhibits robustness issues against these vulnerabilities, as detailed in § 3.3.
Note that the modular evaluation relies on the data generation method. The data generation method produces contrastive data clusters with the same query logic, ensuring a fair evaluation (See Figure 1). Another key aspect of the design choice is the significant combinatorial potential inherent in the proposed data generation process, influenced by various factors. Specifically, a manually crafted database with a modest number of attributes and rows suffices for testing purposes, given the extensive variety of query permutations made possible through the combinatorial approach. A comprehensive analysis of these concepts is presented in Appendix B.3 and C.
Algorithm 1 Grounded Data Generation Process.
1: SQL Template Generator $g_{\mathrm{sql}}$, Text Template Generator $g_{\mathrm{t}}$, Semantic criteria for SQL template generation $\mathcal{C}_{\mathrm{sql}}$, Linguistic criteria for text template generation $\mathcal{C}_{\mathrm{t}}$, Database $\mathcal{D}$, Database Schema $\mathcal{S}=\left\{S_{1},S_{2},\ldots,S_{n}\right\}$, where each $S_{i}$ is a schema for a table
2:$Q\leftarrow\emptyset$triangleright Used to store final evaluation data
3:$\mathcal{S}_{\text{target}}\in\mathcal{S}$triangleright Step 1: Define schema of a domain (§3.1)
4:$\{tpl_{\text{sql}}\}\leftarrow g_{\mathrm{sql}}(\mathcal{S}_{\text{target}},\mathcal{C}_{\mathrm{sql}})$triangleright Step 2: Generate SQL templates applying semantic criteria (§3.1)
5:for $tpl_{\mathrm{sql}}\in\{tpl_{\text{sql}}\}$ do
6:     $\{tpl_{\text{t}}\}\leftarrow g_{\mathrm{t}}(tpl_{\mathrm{sql}},\mathcal{C}_{\mathrm{t}})$triangleright Step 3: Generate text templates applying linguistic criteria (§3.1))
7:     $P\leftarrow$ extract placeholders from $tpl_{\mathrm{sql}}$
8:     for each $p\in P$ do
9:         $C\leftarrow$ extract column from $p$
10:         $V_{p}\leftarrow$ query $\mathcal{D}$ for distinct values of $C$
11:     end for
12:     $Comb\leftarrow$ Cartesian product of $\{V_{p}:p\in P\}$
13:     for each $comb\in Comb$ do
14:         $q_{\mathrm{sql}}\leftarrow$ substitute $comb$ into $tpl_{\mathrm{sql}}$triangleright Step 4: Generate SQL queries (§3.2)
15:         $a\leftarrow$ query $\mathcal{D}(q_{\mathrm{sql}})$triangleright Step 5: Query DB for answers (§3.2)
16:         $\{q_{t}\}\leftarrow$ substitute $comb$ into $tpl_{\mathrm{t}}\in\{tpl_{\text{t}}\}$triangleright Step 6: List of text queries (§3.2)
17:         $Q\leftarrow Q\cup\{(\{q_{t}\},a)\}$
18:     end for
19:end for
20:return $Q$
3.1 Generating Query Templates
This process begins by taking a database schema as input and subsequently produces SQL query templates along with their corresponding textual representations.
Database Schema
Database schema serves as the blueprint that defines how classes of entities are organized and the relations among them. Appendix A demonstrates the reasons why we use database schema for knowledge representation. Let a database schema with a collection of $N$ tables $S=\left\{S_{1},S_{2},\ldots,S_{N}\right\}$, where each $S_{i}$ is a schema for a table. A table schema $S_{i}$ consists of the table name $T_{i}$, a set of attributes $\left\{A_{1},A_{2},\ldots,A_{M}\right\}$ and a set of constraints, i.e., $S_{i}=\left(T_{i},\left\{A_{1},A_{2},\ldots,A_{m}\right\}\right.,C)$, where the constraint $C$ demonstrates the primary key $PK$ and foreign keys $FK$. Specifically, $FK\left(T_{i}.A_{k}\rightarrow T_{j}.A_{h}\right)$ indicates a foreign key $A_{k}$ in table $T_{i}$ referencing the primary key $A_{h}$ in table $T_{j}$. The schema defines the knowledge structure of an entity, which can be enough to generate query logic that is not tied to any specific linguistic expressions and sensitive data.
Generating SQL Templates
[…]
Generating Text Templates
$tpl_{\mathrm{sql}}$ can be instantiated into various text templates, each denoted as $tpl_{t}$. This process, driven by the text template generator $g_{\mathrm{t}}$, leverages linguistic criteria $\mathcal{C}_{\mathrm{t}}$ to transform $tpl_{\mathrm{sql}}$ into natural language forms that align with specified linguistic characteristics, e.g., complexity, length and stylistic nuances. The text template generator $g_{\mathrm{t}}$ is also operationalized using a sophisticated language model, such as GPT-4, which is adept at producing diverse linguistic variations of the same semantic content.
$\{tpl_{\text{t}}\}=g_{\text{t}}(tpl_{\text{sql}},\mathcal{C}_{\text{t}})$
, where $\{tpl_{\text{t}}\}$ signifies the resulting list of text templates.
3.2 From Templates To Evaluation Data
Utilizing a database $\mathcal{D}$ that aligns with the predefined schema, the framework samples a diverse range of queries along with their corresponding ground-truth answers. This dataset, derived from SQL and text templates filled with database content, provides a rich source for evaluation.
Generating Queries Via Placeholder Fill-in
The placeholder fill-in step populates the templates with actual data from the rows of tables in $\mathcal{D}$. Specifically, the approach employs a SELECT query format: “SELECT DISTINCT {column_name} FROM {table_name};” to ensure variety in the data points. It handles SQL query templates with multiple placeholders by employing combinations of placeholders and using Cartesian products to generate multiple query permutations. This approach leads to a comprehensive set of SQL queries ($q_{\mathrm{sql}}$) and their natural language equivalents, the text queries ($q_{t}$).
Generating Answers
Answer generation is integral to completing our dataset. Each SQL query ($q_{\mathrm{sql}}$) is executed against $\mathcal{D}$ to match with a factual answer $a$. Each answer is then paired with the relevant text queries $\{q_{t}\}$, forming the basis of our evaluation data.
Figure 2: Scaled Generation of Query-answer Pairs With Combinatorial Power. The step 3 and step 4 are independent of each other and depend only on the SQL templates.
3.3 Modular Evaluation
Assessing an RAG model’s overall performance may obscure insights into the robustness or weaknesses of its constituent modules (refer to Section 4.2). It is essential, therefore, to evaluate the contributions of the retrieval and language modeling (LM) components individually.
We introduce three group types to distinguish the knowledge gap and the robustness issue, as shown in Figure 1. All textual queries $Q$ generated from a specific SQL logic $q_{\mathrm{sql}}$ can be grouped semantically. [^1] Refer to Appendix D for an illustration. Depending on the model $M$’s performance, we categorize $q_{\mathrm{sql}}$ (and consequently, the group $Q$) into three principal categories:
  • Gap Groups emerge when the model $M$ consistently provides incorrect answers for every text query within a subset $S\subseteq Q$ produced through the data generation process. This reflects a deficiency in $M$ ’s knowledge or capability. Formally, this scenario is described as:
    $\forall q_{t}\in S,\neg M\left(q_{t}\right)$
    wherein $\neg M\left(q_{t}\right)$ signifies an inaccurate response from $M$ to the text query $q_{t}$.
  • Robust Groups are established when model $M$ accurately responds to all text queries within the subset $S$, evidencing a robust comprehension of the SQL logic. Formally, this is articulated as:
    $\forall q_{t}\in S,M(q_{t})$
  • Non-robust Groups are characterized by model $M$’s exhibited capability to potentially address the query logic, yet there exists at least one query $q_{t}$ within $S$ that $M$ inaccurately predicts. This is formally denoted as:
    $\left(\exists q_{t}\in S,M(q_{t})\right)\land\left(\exists q_{t}\in S,\neg M(q_{t})\right)$
Defective Modules Responsible for The Robustness Issue
Gap groups mostly reflect the limited coverage of the retrieval database. The analysis of the semantics of queries in gap groups can help expand the database. However, defective modules responsible for the robustness issue need to be further identified by answering the following questions: Does the retrieval module source and provide the sufficient context needed for a given query? With the sufficient context at its disposal, is the LM capable of drawing accurate inference? Utilizing a “context comparison” methodology enables pinpointing the LM’s role in generating erroneous predictions when adequate context is provided. By contrasting the document indices $\text{idx}_{c}$ of each non-robust instance against indices from accurate predictions within analogous semantic groups against those linked to correct predictions within the same semantic groups $I=\text{idx}_{1},\text{idx}_{2},\dots$, we evaluate the sufficiency of context. This assumes the LM’s reliance on external information for query responses, mirroring our scenario and prevalent industrial applications. Should any index in $I$ match $\text{idx}_{c}$, we reclassify the outcome as a successful retrieval yet an LM shortfall. This strategy allows for precise identification of LM-related non-robust instances. Nevertheless, the absence of such matches does not conclusively indicate retrieval failures. Given our experimental setup with a synthetic database comprising 10 documents, this “context comparison” method effectively identifies all document combinations across a moderate query range (3 to 6) within semantically related groups. This framework lays the groundwork for future explorations into more intricate cases.
A reliable robustness metric is necessary to allow precise identification of detective modules. We quantify robustness as the fraction of instances correctly answered, excluding those in gap groups:
$\begin{split}R=\frac{\text{\# with Correct Predictions}}{\text{Total \#}-\text{Total \# in Gap Groups}}\end{split}$ (1)
, where “#” represents “the number of instances”. By omitting gap groups, this approach ensures that the metric reflects true robustness, distinct from knowledge gaps. Consequently, accuracy Acc is articulated as:
$\begin{split}\text{Acc}=R\times\frac{\text{Total \#}-\text{Total \# in Gap Groups}}{\text{Total \#}}\end{split}$ (2)
This formulation reveals how the proportion of instances within gap groups $\lambda$ influences the efficacy of accuracy as a measure of robustness. Specifically, a higher $\lambda$ value in the robust compared to the non-robust cluster may obscure the model’s non-robust characteristics. Section 4 will empirically demonstrate the crucial importance of excluding gap groups for accurately assessing system vulnerabilities.
4 Case Study: Question Answering on Industrial Projects
This section presents a case study focusing on question answering about industrial projects. It serves to critically evaluate both existing reference-free evaluation frameworks and the newly proposed evaluation framework
System Overview
Our study incorporates two distinct systems: 1) Dense Retrieval + GPT3.5  [^2] OpenAI’s gpt-3.5-turbo-16k: It is deployed in a commercial context with secure access to the retrieval module via the Qdrant API, ensuring data privacy. The system employs an OpenAI embedding model for document embedding, providing a basis to assess the reference-free evaluation framework’s reliability (§4.1). 2) Simple Sparse Retrieval + GPT3.5: The absence of comprehensive research into the assured robustness of dense retrieval systems makes it difficult to pinpoint distinct robust or vulnerable characteristics with confidence. Consequently, we employ a straightforward keyword-matching retrieval approach. Although this method is less prevalent in real-world applications, this method’s inherent shortfall in handling lengthy queries makes it exceptionally ideal to validate our evaluation framework in conducting reliable hypothesis testing (§4.2).
Synthetic Knowledge Base
Concurrently, a synthetic dataset has been created, consisting of project documents from the fictitious company, Aurp, serving as a retrieval database. It facilitates a precise evaluation of the retrieval component and the language model, since the knowledge contained within the synthetic data is not present in the training datasets of large language models. In scenarios where retrieval fails to procure adequate context, the language models are definitely unable to accurately respond to queries. Additional information on this methodology can be found in Appendix F.
Mapping Domain Knowledge into Database Schema
Our approach distills domain knowledge into a database schema centered around three pivotal entity types in project management: Project, Client, and Employee.  [^3] This triad are reflective of the core components outlined in the Project Management Body of Knowledge (PMBOK), a widely recognized standard in the field of project management (Guide, 2001).This schema encapsulates the quintessential entity types found within a corporate setting. With the scalability of our framework, the limited amounts of entity types and associated attributes and entities lead to large amounts of instances for evaluation: 198 examples for the commercial case (with totally 12 attributes and 11 rows) and 942 examples for the synthetic scenario (with 14 attributes and 49 rows in total). The details of the database schema and the amounts of columns and rows for each table are shown in Appendix E and G), respectively.
4.1 Is Reference-free Evaluation Reliable?
Leveraging ground truths, our evaluation framework sets a benchmark for reference-free assessments. This section delves into the reliability of reference-free evaluation methods, utilizing examples generated for the commercial case. The focus is to discern the reliability of such evaluation approaches: How trustworthy are these reference-free protocols? From what perspectives can they be relied upon in domain-specific contexts?
Reference-free Evaluation Protocols
We assess two reference-free evaluation protocols: SelfCheck and RAGAS. RAGAS-Fact utilizes the context-query-response triplets to assess the veracity of responses. It evaluates the faithfulness of a response by calculating the ratio of claims grounded on the context to the total claims made. This process, termed as Ragas-Fact, involves identifying statements that hold atomic facts, following the methodology outlined by Chern et al. (2023). SelfCheck, as introduced by Manakul et al. (2023), relies on stochastically generating responses from the model under the premise that incorrect answers are unlikely to repeatedly yield identical outcomes through random sampling. This principle, initially applied to LLMs, is adapted for our analysis of RAG systems. In a departure from its original application, we enhance the evaluation prompt to include not only stochastic responses but also the query itself. Refer to Appendix H for details. To generate four stochastic samples, we adjust the temperature setting to 1.0, contrasting with a temperature of 0.0 used to procure the primary response. A response is deemed correct if it aligns consistently across all stochastic samples.
RAGAS’ Reliable Assessment of Correct Predictions Amidst Broader Unreliability in Reference-Free Evaluation
To gauge reliability, we devise two metrics: the ratio of predictions judged as accurate among all those classified as correct by the evaluation, and the ratio of predictions deemed accurate within the set of instances that were indeed correctly predicted. When considering reference-free evaluation as a binary classification task, these metrics align with precision and recall, respectively. In this context, the positive class represents correct predictions, and the negative class represents incorrect ones. The ground truth corresponds to the actual accuracy of model predictions, while the evaluation task is to determine the correctness of these predictions. As shown in Table 1, RAGAS demonstrates effectiveness in assessing accurately predicted instances, yet it achieves only a 19% accuracy rate across all evaluations, highlighting a deficiency in correctly identifying erroneous predictions. This trend suggests a propensity for inflating model performance assessments, likely due to the RAG system producing responses that, while contextually relevant, fail to directly address the intended query. SelfCheck performs even worse than RAGAS in both cases. SelfCheck’s performance falls short of RAGAS’s, underperforming in both precision and recall. The preliminary insight suggests that reference-free evaluation methods are inherently prone to inaccuracies, underscoring the critical need for a robust, reference-based evaluation framework to ensure dependable model assessment.
Precision Recall
Ours (Upper Bound) 1 1
RAGAS-Fact 19% 92%
SelfCheck 15% 50%
Table 1: Reliability of reference-free evaluation protocols.
Robust Non-robust Success?
Accuracy Baseline: Unbalanced Sampling (US) 0.31 0.43 No
US+Modular Evaluation (Retrieval) 0.36 0.46 No
Stratified Sampling (SS) 0.31 0.29 Yes
SS+Modular Evaluation (Retrieval) 0.38 0.35 Yes
Robustness Baseline: Unbalanced Sampling (US) 0.84 0.85 No
US+Modular Evaluation (Retrieval) 0.99 0.93 Yes
Stratified Sampling (SS) 0.78 0.74 Yes
SS+Modular Evaluation (Retrieval) 0.94 0.87 Yes
Table 2: Evaluating the efficacy of accuracy and designed robustness metrics in distinguishing between robust and non-robust clusters. Modifications include retrieval-specific adjustments by excluding LM-induced non-robust examples.
4.2 Empirical Validation of GRAMMAR
This section will validate design choices of identifying defective modules in the proposed framework.
The Importance of Modular Evaluation: Avoiding The Masking Effect of Comprehensive Assessments
Analysis presented in Table 2 underscores the value of segmenting the evaluation to specifically examine the retrieval process. The non-robust cluster actually receives higher scores (better performance) than the robust cluster regarding both the accuracy and robustness metrics. However, under the modular evaluation, the evaluation result correctly reflects the non-robust cluster. This underscores the importance of evaluating the retrieval function independently; failing to do so may allow language model inaccuracies to obscure the true performance metrics, potentially masking retrieval robustness issues. Although instances causing such distortions were not dominant in our dataset, they highlight a critical evaluation nuance.
The Importance of Excluding Gap Groups for Robustness Measurment
Highlighting the necessity of omitting gap groups, our ablation study implements two distinct approaches: 1) Comparing the Robustness Metric with Accuracy; 2) Comparative Analysis of Stratified Sampling and Unbalanced Sampling. Stratified sampling maintains a consistent $\lambda$ across both robust and non-robust clusters. By prompting the text template generator, an equal number of text queries are generated for each query logic. To explore the effects of unbalanced sampling, that can occur in the real world, we randomly reduce examples from query groups related to “clients” and “employees”.
Table 2 demonstrates that traditional accuracy metrics falter in distinguishing clusters based on vulnerability attributes, particularly when $\lambda$ varies between the hypothesis-driven cluster and its counterpart (Row 2 under the “Accuracy” category). Stratified sampling ameliorates this issue, enabling a precise identification of both robust and non-robust clusters (Row 4 under the “Accuracy” category). Importantly, the implementation of a specifically designed robustness metric ensures consistent and accurate identification of the robust cluster, irrespective of $\lambda$ variations (Row 2 and 4 under the “Robustness” category). This approach, integrating stratified sampling with our bespoke robustness metric, providing a reliable analysis.
5 Conclusions
In our work, we have developed an evaluation framework tailored for Retrieval-augmented Generation (RAG) systems within an industrial context, enhancing the reliability and interpretability of evaluations. Additionally, we have exposed the limitations of current reference-free evaluation methods, particularly their ineffectiveness in accurately assessing incorrect predictions.
Reproducibility
The entire evaluation framework, encompassing both the code and the prompts for the LLM, will be made open-source and available to the academic and industrial community. This initiative aims to facilitate the wider use and adaptation of our framework for research and real-world implementations. In terms of the systems and data evaluated to enable result replication, we plan to publish the model, including the documents intended for retrieval, as well as all templates and queries developed for the hypothetical Aurp scenario. However, the model designated for commercial use and its associated knowledge base will remain confidential and unpublished to honor non-disclosure agreements and protect proprietary information.
6 Limitations
The challenges related to the presence of multiple correct answers (refer to Appendix B.1) and the constraints in expressiveness posed by database schema and SQL (refer to Appendix B.4) ultimately result in difficulties in generating queries that demand multi-step reasoning and free-form responses. Nonetheless, as demonstrated in the case study, straightforward queries prove to be sufficient for identifying and debugging the intrinsic robustness issues within the models.
Appendix A Knowledge Structure and Representation
Commonly, the knowledge is formalized by entities along with their relations, e.g., a knowledge graph, ontology and a database schema. Instead of using knowledge graphs containing concrete entities and their relations, we use the schema-based definition, where the schema is characterized by entity types (classes of entities). An entity type is defined by its name, its attributes and relations with other entity types that will be concretized as components in a database, i.e., the table name, columns and foreign keys.
The definition of entity types can provide combinatorial expansion of data generation, protecting the leak of private information (normally stored as rows of tables) during the data generation process and the ability of retrieving ground-truths via Structured Query Language (SQL). The three advantages are the reasons why we represent knowledge in database schema rather than knowledge graphs. It disentangles the private data stored in the database with universal metadata expressed in the database schema. Since schemas and templates contain only meta-data for structuring the real data, there is no private issue to use commercial large language models for this process.
    You are a SQL query Template Generator: Generate ACCEPTABLE SQL query     templates with placeholders according to the give data schema and     requirements. A simple example of an acceptable SQL query template is:     SELECT Industry FROM Company WHERE Name = ’[Company.Name]’;     You must follow the basic criteria below except for other requirements:     ##CRITERIA##     - The placeholder format should be a combination of a table name and     a column name, enclosed within square brackets, e.g., ’[User.Name]’.     - Use only ’SELECT’ queries.     - Select specific column(s) instead of using ’*’. Avoid projecting     attributes that appear in the predicate.     - The selected and condition columns in the query MUST BE MEANINGFUL and     DESCRIPTIVE to ensure the queries are easily understood by non-technical     users.     - Avoid using technical column names that don’t clearly signify the nature     of the entities or objects involved, e.g., column for semantically void     record identifiers.     - Do not create redundant or semantically duplicated queries when translated     into natural language.     - Each query must contain at least one parameter placeholder in the WHERE     clause.     - Ensure the query yields a specific and singular answer to avoid multiplicity     issues, thus facilitating accurate chatbot evaluation.      {SPECIFIC_REQUIREMENTS}     - If no acceptable SQL template can be generated with the given table and     column information, do not generate any text.     ##RESPONSE FORMAT##     - Output each SQL template as a single line, without any prefix or suffix.     - Do not include any other text in your response, even something like     ##RESPONSE_END##.     ##DATA SCHEMA##     {GIVEN_SCHEMA}     ##RESPONSE_START##
Table 3: A prompt template designed to guide the LLM in functioning as a controllable SQL template generator
Appendix B Details of SQL Template Generation
B.1 Constraints/Guidelines for Creating SQL Templates
The constraints limit the types of templates that can be generated. This section demonstrates two important constraints, which are verbalized as the prompt for LLMs.
Selecting Attributes That Are Understandable to Humans
As shown in Table 3, the SQL template generator is required to follow the constraint: “The selected and condition columns in the query MUST BE MEANINGFUL and DESCRIPTIVE to ensure the queries are easily understood by non-technical users.”.
Avoiding Answer Multiplicity
In the evaluation of question-answering (QA) models, a unique challenge arises from the existence of multiple valid answers to a single query, which necessitates a nuanced approach to assessing model performance. Consider the question: “Get the name of the client associated with the project named Innovation Precinct” For such a question, a set of correct responses could include any combination of names from a predefined list, such as {Apple, Amazon, Meta, Facebook}. This multiplicity of correct answers underscores the complexity of evaluating QA models, as it requires the assessment mechanism to recognize and validate the full spectrum of possible correct answers rather than comparing the model’s output against a single ’gold standard’ answer. This scenario demands a more flexible method for answer validation that can accommodate the variability in correct responses. Also, evaluation metrics may be required to effectively measure the performance of QA models in handling diverse and equally valid answers. Such metrics must account for the exhaustive set of correct answers and evaluate the model’s ability to retrieve any or all valid responses within the context of the query, thereby ensuring a comprehensive assessment of the model’s understanding and response generation capabilities. Generally, encompassing all valid responses within the database is not feasible. Hence, rather than let the non-robust results affect the accuracy of evaluation, we look for an approach to avoid it, i.e., only generating queries ensuring one ground-truth answer. Specifically, the evaluation of certain SQL queries, such as “SELECT Name, Department, BusinessAddress FROM Employee WHERE JobTitle = ’designer”’, requires a complete and thorough listing of multiple answers that can be dynamically changed. To solve the issue, the solution involves adding a specific criterion to the prompt for generating SQL queries. This criterion, “Ensure the query yields a specific and singular answer”, aims to produce queries that result in a single, clear answer, thereby avoiding the complexities of multiple possible correct answers. For example, with this criterion, the query “SELECT Industry FROM Company WHERE Name = ’[Company.Name]’;” is preferred as it’s likely to yield a singular answer about a company’s industry based on a specific company name. In contrast, without this criterion, queries like “SELECT Name FROM Company WHERE Industry = ’[Company.Industry]’;” are acceptable but may result in multiple names, leading to evaluation difficulties due to data completeness and query multiplicity issues.
B.2 LLM-Based, Human-Aided SQL Generation
We utilize GPT-4 as a SQL template generator for its state-of-the-art capabilities in SQL generation benchmarks (Liu et al., 2023). However, according to our results, LLMs could violate the constraints specified in the prompt. Therefore, the tool incorporates built-in functions that facilitate manual verification by saving, modifying, and reloading SQL templates. As detailed in Table 3, the profile prompt defines its purpose and establishes constraints for generating valid SQL templates.
The following process demonstrates the use of the developed tool to generate SQL templates:
1## a. Automatically generate SQL query templates
2sql_template_generator = SQLTemplateGenerator(connection_string, llm)
3sql_template_generator.set_system_msg()
4sql_templates = sql_template_generator.generate()
5sql_template_generator.save(sql_templates, root_dir, overwrite=True)
The output is saved in "SQLTemplateGenerator/sql_templates.txt" with the following SQL queries:
1SELECT Name FROM company WHERE Sales_in_Billion = ’[company.Sales_in_Billion]’;
2SELECT Industry FROM company WHERE Name = ’[company.Name]’;
3SELECT Headquarters FROM company WHERE Market_Value_in_Billion = ’[company.Market_Value_in_Billion]’;
4SELECT Nationality FROM people WHERE Age = ’[people.Age]’;
5SELECT Graduation_College FROM people WHERE Name = ’[people.Name]’;
6SELECT Year_working FROM employment WHERE Company_ID = ’[employment.Company_ID]’;
7SELECT People_ID FROM employment WHERE Year_working = ’[employment.Year_working]’;
To ensure the reliability of the generated SQL templates, they can be manually verified and modified, and new templates can be added. Modifications and reloading of templates can be performed using the following steps:
1# b. Manually modify SQL query templates (optional)
2
3# c. Load modified SQL query templatesn (optional)
4sql_templates = SQLTemplateGenerator.load(root_dir)
[…]
Below are the modified SQL queries as suggested in the §B.1:
1SELECT Industry FROM company WHERE Name = ’[company.Name]’;
2SELECT Headquarters FROM company WHERE Name = ’[company.Name]’;
3SELECT Nationality FROM people WHERE Name = ’[people.Name]’;
4SELECT Graduation_College FROM people WHERE Name = ’[people.Name]’;
5SELECT e.Year_working FROM employment e JOIN company c ON e.Company_ID = c.Company_ID JOIN people p ON e.People_ID = p.People_ID WHERE c.Name = ’[company.Name]’ AND p.Name = ’[people.Name]’;
B.3 Possibilities Of SQL Templates
One Table/Entity
Given a schema with only one table, the possibilities for SQL templates can be analyzed by considering the basic SQL operators like SELECT, WHERE. Here’s a breakdown:
  • Variation in Selected Attributes (SELECT): The number of SQL templates varies based on the combination of columns selected. If the “Company” table has $n$ columns, then theoretically, there are $12^{n}-$ 1 possible combinations of columns for selection (excluding the case where no column is selected).
  • Conditions in Queries (WHERE): Each SQL query can include zero or more conditions in the WHERE clause. The number of possible conditions is determined by the number of columns, the type of each column (text, numeric, date, etc.), and the range of operators applicable to these types (like =, <, >, LIKE, IN for text columns; =, !=, <, >, BETWEEN for numeric columns). The complexity increases combinatorially with multiple conditions combined using AND/OR.
Two Entities
The relations between tables further amplifies the number of possible templates. For example, with two entities, “Company” and “Project,” and their associative table “Company_Project”, the possibilities for SQL templates expand significantly due to the introduction of joins and more complex WHERE clauses. Let’s break down the possibilities.
  • Selection Variations (SELECT): The number of SQL templates grows with the combination of columns selected across the three tables: “Company”, “Project” and “Company_Project”. If “Company” has n columns, “Project” has m columns, and “Company_Project” has p columns, the possible combinations for selection are $(2^{n}-1)\times(2^{m}-1)\times(2^{p}-1)$.
  • Join Conditions (JOIN): The introduction of the associative table “Company_Project” allows for meaningful JOIN operations between “Company” and “Project.” Templates can include joins like Company JOIN Company_Project ON condition and Project JOIN Company_Project ON condition, or a multi-table join linking all three. The variety of JOIN conditions adds another layer of complexity to the possible templates.
  • WHERE Clause Complexity: With more tables, the WHERE clause can include a wider range of conditions, potentially involving attributes from any of the three tables. The complexity increases with the number of columns and their types across all tables, and combinations of these conditions.
B.4 Limitations
Scope of Using Automated SQL Generators
The effectiveness of queries generated by LLMs hinges on the meaningfulness of table and column names to accurately reflect the essence of the entities or concepts they represent. If these names lack contextual clarity, the resulting queries may be impractical. To mitigate this, practitioners can conduct an informal assessment or consult established benchmarks that evaluate LLMs’ proficiency in domain knowledge and commonsense reasoning, for instance, (Zhong et al., 2023; Hendrycks et al., 2021).
Queries Beyond SQL Expressiveness
While SQL and relational algebra offer a wide range of operations enabling the formulation of numerous user queries, certain semantic nuances exceed the expressive capabilities of SQL. The following examples illustrate such limitations:
  • Real-world queries often contain ambiguity and subjective interpretations that SQL struggles to accommodate. For instance, the term “major” in the query “What are some major rail projects we’ve been involved with?” does not directly translate into SQL criteria without additional interpretative steps.
  • SQL queries are typically structured to elicit specific, predefined responses, contrasting with the open-ended nature of many real-world inquiries that seek exploratory or comprehensive answers.
Appendix C Breakdown of Scalable Data Generation
The proposed data generation process of generating queries via LLM, which includes both the creation of templates and the formulation of queries, can be succinctly represented by the formula (see Figure 2 for a visual representation):
Total Query $\text{Total Query Variations}=M\times N\times Q,$ (3)
where “M” denotes the array of SQL templates that can be extracted from a given schema, “N” refers to the vast number of text templates that can be produced from a single SQL template, showcasing the flexibility of natural language, and “Q” accounts for the broad spectrum of text queries that can be generated from a single text template, with this diversity arising from different combinations of unique values filling the placeholders.
Let’s break down each step of the proposed data generation process:
  • 1 schema $\Rightarrow$ M SQL Templates: The number of possible SQL templates (M) that can be generated from a defined schema is influenced by the SQL operators used, the number of tables, and the columns in the database schema. Each combination of tables and columns, along with different SQL operators (like SELECT, WHERE, JOIN), can lead to a unique SQL query template, e.g., the exponential complexity of the SELECT operator, and the combinatorial increase of possible predicates (See Appendix B.3 for details). Note that while the theoretical maximum is high, practical and meaningful queries will be a smaller subset, as specified in the case study and Appendix B.
  • 1 SQL Template $\Rightarrow$ N Text Templates: For each SQL template, there can be an arbitrary number (N) of textual expressions. This variation arises from the different ways to linguistically express the same SQL query due to the flexibility and richness of natural language, e.g., the variation of synonyms and sentence structures.
  • 1 Text Template $\Rightarrow$ Q Text Queries: Each text template can lead to a number of text queries (Q), depending on the unique values available for each placeholder. With more than one placeholder, the potential for growth in the number of text queries is combinatorial, barring semantic conflicts (context conflict), where certain combinations of column values may not be semantically valid. Overall, this framework demonstrates a combinatorial expansion at each transition stage, especially notable in steps involving natural language due to its inherent variability. However, this growth is tempered by practical constraints such as the meaningfulness of queries (semantic validity) and the actual data distribution in the database. The exponential increase is most pronounced in the transition from text templates to text queries, where the permutations of placeholders can lead to a vast array of unique query possibilities.
Appendix D Semantic Group
Textual queries that originate from a specific SQL logic $q_{\text{sql}}$ are organized into a semantic group because of their underlying shared semantics. For instance, the SQL query “SELECT C.address FROM COMPANY C WHERE C.CompanyName = ’Carlton Innovation Precinct’“ results in the creation of a semantic group $Q=\{q_{t1},q_{t2},\ldots\}$, with each $q_{t}$ representing a distinct textual interpretation. Examples include $q_{t1}$ as “What is the address of ’Carlton Innovation Precinct’?” and $q_{t2}$ as “Where is ’Carlton Innovation Precinct’ located?”.
Appendix E Database Schemas Overview
Figures 3(a) and  3(b) illustrate the database schemas used for an actual industrial context and a fabricated scenario, respectively, within our research.
(a) Aurecon.
(b) Aurp.
Figure 3: Entity-Relationship Diagrams for Data Generation
Attribute # Template # Row # Queries #
SQL SQL (Extra Constraints) Text SQL Text
Project 5 7 3 18 4 13 78
Employee 5 12 5 30 3 14 84
Client 2 6 2 12 4 6 36
Total / / / / / 33 198
(a) Aurecon
Attribute # Template # Row # Queries #
SQL Text SQL Text
Project 6 5 15 10 50 300
Employee 5 4 12 29 87 522
Client 3 2 6 10 20 120
Total / / / / 157 942
(b) Aurp
Table 4: Summary of DB Rows and Columns and Templates/Queries
Appendix F Synthetic Knowledge Base for Evaluating Retrieval-Augmented Language Models
The process of generating synthetic data for a synthetic environment involves creating comprehensive, detailed, and interconnected datasets for a fictitious company named Aurp, which specializes in structural engineering consultancy services. This multi-step procedure includes:
  1. 1.
    Generating Company Profile: Initially, a company profile for Aurp is created, detailing its foundation year, headquarters location, CEO, number of employees, and the services it offers, such as bespoke architectural solutions, sustainable urban planning, and structural health monitoring.
  2. 2.
    Generating Organizational Structure: Next, a project-oriented organizational structure is established, naming key positions and employees within the company, similar to real-world firms. This includes a wide range of roles from executive positions to specialized engineers and support staff.
  3. 3.
    Generating Employee Information: For each employee listed in the organizational structure, detailed job titles, departments, and direct supervisors or managers are fabricated, creating a network of relationships and reporting lines within the company.
  4. 4.
    Generating Client Information: The process creates a list of ten clients across various industries—ranging from technology and real estate to hospitality and healthcare—each with specified locations, thereby illustrating the company’s diverse portfolio.
  5. 5.
    Generating Projects: Specific projects are then devised, including names, locations, start and end dates, clients, project directors, and project managers. This step integrates previously generated data (client and employee information) to create realistic project scenarios.
  6. 6.
    Generating Project Reports: This step involves synthesizing data from the previous steps to produce detailed analyses, updates, and outcomes of the various projects, which may or may not contain detailed information for clients and employees.
This elaborate and interconnected data creation method is tailored for evaluating retrieval-augmented language models. It serves as a critical tool for exploring how these models handle data that is entirely fictional and thus not present in their training datasets. The synthetic nature of this environment provides a unique testbed for separating and assessing the distinct capabilities of the retrieval component and the language model component in isolation. Because the knowledge in the synthetic data does not appear in the training data of large language models.
Appendix G Assessing Scalability with Limited Entity Types, Rows, and Attributes
Our scalability assessment focused exclusively on the utilization of the SELECT operator and the WHERE clause for each table schema. As illustrated in Table 4(a) and 4(b), imposing constraints to generate relevant queries significantly reduces the potential quantity of SQL templates based on attribute count. Despite these restrictions, a minimal set of attributes and rows still yields hundreds of viable query-answer pairs.
Appendix H Prompt Templates for LLM-Based Evaluation
For SelfCheck, given that our responses are typically concise, we opt for assessing the overall response rather than rather than performing a sentence-by-sentence validation as in the original implementation, as shown in Table 5.
SelfCheck from (Manakul et al., 2023)
Context: {context}
Sentence: {sentence}
Is the sentence supported by the context above?
Answer Yes or No: SelfCheck-QA
Query: {query}
Answer A: {answer}
Answer B: {stochastic_answer}
Do both answers address the query with equivalent meaning?
Use only “Yes” or “No” for your evaluation:
Ragas from (Es et al., 2023)
Natural language inference. Use only ’Yes’ (1), ’No’ (0) and ’Null’ (-1) as verdict.
context: {context}
statement: {statement}
verdict:
Ours
Evaluate the accuracy of the given response in relation to the true answer for the specified query. After evaluating, provide a judgement as either “Correct” or “Incorrect” based on whether the ##Given Response## accurately matches the ##True Answer##.
##Query##: {query}
##True Answer##: {true_answer}
##Given Response##: {given_response}
##Judgement##:
Table 5: Prompt templates for LLM-based Evaluation.